T- bill rate is 4%. A risk-averse investor with a degree of risk aversion A = 3 invests entirely in a risky portfolio with a standard deviation of 24%. What should the risky portfolio's expected return be?
Answer #1

Expected Return[E(r)] - Rf = 0.5 * A * (Standard Deviation)2

Expected Return[E(r)] = 0.5 * 3 * (0.24)2 + 0.04

= 0.0864 + 0.04

Expected Return[E(r)] = 0.1264 or 12.64%